In an experiment to determine the period of oscillation of a loaded spiral spring, the equation of a simple harmonic oscillator a=−(k/m)x, where the symbols have their usual meaning, was used. Write down the equation for the angular frequency ω in terms of k and m, the effective mass of the system
1
Expert's answer
2017-06-29T11:13:06-0400
The equation of a simple harmonic oscillator is
a=-ω^2 x.
Thus,
ω^2=k/m.
The equation for the angular frequency ω in terms of k and m, the effective mass of the system is
ω=√(k/m).